OpenStudy (anonymous):

A moon orbiting a planet in the solar system has an orbital period of 5.0 x 10^4 s, and the distance between the center of the moon and the planet is 2.0 x 10^8 m. What planet is it? a. Mars (mass = 6.42 x 10^23 kg) c. Venus (mass = 4.87 x 10^24 kg) b. Neptune (mass = 1.02 x 10^26 kg) d. Jupiter (mass = 1.90 x 10^27 kg)

OpenStudy (compphysgeek):

to solve this problem you use the expression for centripetal force \[ F_C = m \omega^2 r \] and set it equal to the gravitational force \[F_G = G \frac{m_P m_M}{r^2}\] with \(m_P\) the mass of the planet that you're looking for, \(m_M\) the mass of the moon, \(r\) the distance between moon and planet, \(\omega = \frac{2 \pi}{t}\) and the gravitational constant \(G = 6.67 \times 10^{-11} [G]\). You then solve for the planet's mass. \[ m_M = \frac{\omega^2 r^3}{G}\]

OpenStudy (compphysgeek):

the last one is of course \(m_P = \)

OpenStudy (anonymous):

was he right

OpenStudy (compphysgeek):

I was right ;)